# Continuous Input Embedding Size Search For Recommender Systems

### v4

@@ Abstract.
Latent factor models are the most popular backbones for today’s recommender systems owing to their prominent performance. Latent factor models represent users and items as real-valued embedding vectors for pairwise similarity computation, and all embeddings are traditionally restricted to a uniform size that is relatively large (e.g., 256-dimensional). With the exponentially expanding user base and item catalog in contemporary e commerce, this design is admittedly becoming memory-inefficient. To facilitate lightweight recommendation, reinforcement learning (RL) has recently opened up opportunities for identifying varying embedding sizes for different users/items. However, challenged by search efficiency and learning an optimal RL policy, existing RL-based methods are restricted to highly discrete, predefined embedding size choices. This leads to a largely overlooked potential of introducing finer granularity into embedding sizes to obtain better recommendation effectiveness under a given memory budget. In this paper, we propose continuous input embedding size search (CIESS), a novel RL-based method that operates on a continuous search space with arbitrary embedding sizes to choose from. In CIESS, we further present an innovative random walk-based exploration strategy to allow the RL policy to efficiently explore more candidate embedding sizes and converge to a better decision. CIESS is also model-agnostic and hence generalizable to a variety of latent factor recommender systems, whilst experiments on two real-world datasets have shown state-of-the-art performance of CIESS under different memory budgets when paired with three popular recommendation models.
@@ 1. Introduction
The root cause is that, each user/item embedding in the conventional embedding table shares the same fixed embedding size. An example from (Lian et al., 2020) shows that a recommender model embedding 10 million items into 256-dimensional vectors can exceed 9 GB memory consumption in a double-precision float system. As such, researchers have developed better solutions to compress the embedding table while maintaining its expressiveness. The most primitive recommender uses binary codes as an alternative to embeddings (Zhang et al., 2016), which was soon supplanted by another line of methods that compress fixed-sized embeddings into a lightly parameterized component. For example, codebooks have been proposed to store the embedding latent vectors efficiently in (Lian et al., 2020; Shi et al., 2020; Xia et al., 2023). (Xia et al., 2022) introduced the semi-tensor product operation to tensor-train decomposition to derive an ultra-compact embedding table, and (Kang et al., 2021) encoded users and items into hash codes and applied neural layers to learn their dense representations. However, these methods must be redesigned and retraining for different memory budgets to maintain optimal recommendation performance.
@@ 2. Methodology
CIESS has two main components that work alternately during training: (1) a recommendation model $F_{\Theta}(\cdot)$ parameterized by $\Theta$; and (2) the RL-based search function $G_{\Phi}(\cdot)$ parameterized by $\Phi$. A schematic view of CIESS’s workflow is shown in Figure 1. In each optimization iteration of CIESS, the recommender $F_{\Theta}(\cdot)$ adjusts its user/item embedding sizes to the ones provided by the policy $G_{\Phi}(\cdot)$, then updates its parameters $\Theta$ with training samples. Afterwards, $F_{\Theta}(\cdot)$ is evaluated on a hold-out dataset, where the top-$k$ recommendation quality can be measured by common metrics such as Recall at Rank $k$ (Recall@$k$) and Normalized Discounted Cumulative Gain at Rank $k$ (NDCG@$k$) (He et al., 2020). Based on the recommendation quality, the search function $G_{\Phi}(\cdot)$ will be revised, and then updates its embedding size selection for each user/item for the next iteration. In what follows, we unfold the design of CIESS.
@@ 2.1. Base Recommender with Masked Embeddings
Let $\mathcal{U}$ and $\mathcal{V}$ be a set of users $u$ and items $v$, respectively. Their embedding vectors are stored in a real-valued embedding table $\mathbf{E}$ with the dimensionality of $(|\mathcal{U}|+|\mathcal{V}|)\times d_{max}$. It can be viewed as the vertical concatenation of all user and item embeddings $[\mathbf{e}_{u_{1}};...;\mathbf{e}_{u_{|\mathcal{U}|}};\mathbf{e}_{v_{1}};...;\mathbf{e}_{v_{|\mathcal{V}|}}]$, where $d_{max}$ is the initial embedding dimension of all users/items in the full-size embedding table. In other words, $d_{max}$ is also the maximum embedding size in the search space.
@@
With the mask $\mathbf{M}$, for each user/item, we can retain the first $d_{n}$ elements of its full embedding while setting all succeeding dimensions to 00. It is worth noting that, performing embedding sparsification by masking unwanted dimensions with zeros is a commonly adopted approach in lightweight recommender systems (Liu et al., 2021; Lyu et al., 2022), as the resulting embedding table can take advantage of the latest sparse matrix storage techniques (Virtanen et al., 2020; Sedaghati et al., 2015) that bring negligible cost for storing zero-valued entries.
@@
where our target sparsity ratio $c$ ($10<c<$ 1) specifies the percentage of parameters to be pruned (i.e., zeroed out in our case) in the final sparse embedding matrix $\mathbf{E}\odot\mathbf{M}$.
@@ 2.2.2. State
where $f_{n}$ is the popularity of the user/item, normalized by the corresponding maximum/minimum frequency among all users $f_{max}=\max_{u\in\mathcal{U}}(f_{u})$/$f_{min}=\min_{u\in\mathcal{U}}(f_{u})$ or items $f_{max}=\max_{v\in\mathcal{V}}(f_{v})$/$f_{min}=\min_{v\in\mathcal{V}}(f_{v})$, which are observed from training data. $d_{n}$ is the current embedding size allocated to a user/item, and $q_{n}$ quantifies the changes in the recommendation quality when the embedding size decreases from $d_{max}$ to the current $d_{n}$. Compared with (Zhao et al., 2021b), incorporating this quality indicator into the state is able to help trace the impact from the most recent action (i.e., embedding size) to the recommendation effectiveness, which can encourage the policy network to better balance the memory cost and performance with the embedding sizes selected from a vast, continuous action space.
@@
where the choices of $k$ in our paper are $\mathcal{K}=\{5,10,20\}$, and Recall@$k_{u}$ and NDCG@$k_{u}$ denote the evaluation scores w.r.t. user $u\in\mathcal{U}$.
@@ 2.2.3. Reward
The reward $r$ is the feedback to the current policy to guide subsequent embedding size adjustments. Given our goal stated in Eq.(5), the reward should reflect a blend of recommendation quality and space complexity. Therefore, we design the following pointwise reward for a user/item on their current embedding size $d_{u}$/$d_{v}$ selected:
@@ 2.2.4. Action
At every iteration $i$, the policy network in the search function $G_{\Phi}$ predicts an action, i.e., dimension $d_{n}^{i}\in\mathbb{N}_{>0}$ from interval $[1,d_{max}]$ given a user/item state $s_{n}^{i}$. The action is the embedding size for the corresponding user/item. The recommender $F_{\Theta}(\cdot)$ takes this action by altering the embedding sizes of users/items (i.e., updating $\mathbf{M}$ in practice), and yields a reward $r^{i}_{n}$ along with a subsequent state $s_{n}^{i+1}$. The tuple $(s^{i}_{n},d^{i}_{n},r^{i}_{n},s^{i+1}_{n})$ is defined as a transition, which are stored in a replay buffer $\mathcal{B}$ for subsequent training.
@@ 2.2.5. Actor and Critic
Then, starting from $\hat{d}^{i}_{n}$, we perform random walk (with replacement) to obtain a series of actions $\mathcal{Z}_{\hat{d}^{i}_{n}}$, where we specify $|\mathcal{Z}_{\hat{d}^{i}_{n}}|$ to be relatively small to balance efficiency and exploration thoroughness. After obtaining $\mathcal{Z}_{\hat{d}^{i}_{n}}$, we evaluate each action in $\mathcal{Z}_{\hat{d}^{i}_{n}}$ with the critic network $Q_{\mathcal{U}}(\cdot)$/$Q_{\mathcal{V}}(\cdot)$ and greedily select the final action $d^{i}_{n}$ for iteration $i$:
@@ 2.3. Selective Retraining for Sparsified Embeddings
6:         /* Each iteration applies to all $n\in\mathcal{U}\cup\mathcal{V}$/
@@
[…]
@@ 4. Related Work
Latent Factor Recommenders. Neural networks have exhibited superior ability in solving recommendation tasks. Since MLP can learn the non-linear interactions between users and items, feature representation learning with MLP has been commonplace. He et al. (He et al., 2017) proposed Neural Collaborative Filtering consisting of an MLP component that learns the non-linear user-item interactions and a generalized matrix factorization component that generalizes the traditional matrix factorization. Cheng et al. (Cheng et al., 2016) proposed Wide&Deep connecting a wide linear model with a deep neural network for the benefits of memorization and generalization. Apart from MLP-based models, graph-based methods (Wang et al., 2019; He et al., 2020; Yu et al., 2022) have also demonstrated promising capability. Wang et al. (Wang et al., 2019) proposed the graph neural network-based model NGCF to model the user-item interactions with a bipartite graph and propagate user-item embeddings on the embedding propagation layers. Similarly, He et al. (He et al., 2020) resorted to Graph Convolution Network (GCN) and proposed LightGCN that simplified NGCF by including only its neighborhood aggregation component for collaborative filtering. LightGCN performs neighborhood filtering by computing the weighted sum of the user and item embeddings of each layer. Another promising direction, Factorization Machine (Rendle, 2010), along with its deep variants QNFM (Chen et al., 2021a), DeepFM (Guo et al., 2017) and xLightFM (Jiang et al., 2021) have been studied. (Blondel et al., 2016) was also created to mine higher-order interactions. In addition, heterogeneous data such as textual (Gong and Zhang, 2016; Li et al., 2017; Chen et al., 2020) and visual (Wang et al., 2017; Lee et al., 2018) data has also been exploited by several CNN or RNN-based models. Most of these recommenders utilize vectorized embeddings, which can be optimized by CIESS for memory-efficient embeddings.
@@
AutoML for Recommendation. Designing deep recommender models heavily relies on the expertise of professionals. To alleviate the need for human engagement, Automated Machine Learning for Recommender Systems has been studied to automate the design process for recommender models in a task-specific and data-driven manner. So far several research directions including feature selection search (Wang et al., 2022; Luo et al., 2019), embedding dimension search (Liu et al., 2021, 2020; Zhao et al., 2021b; Qu et al., 2022), feature interaction search (Su et al., 2021, 2022), model architecture search (Song et al., 2020; Cheng et al., 2022), and other components search (Zhao et al., 2021a; Meng et al., 2021) have been proposed (Zheng et al., 2022). The first kind reduces computation cost by filtering feature fields based on learnable importance scores (Wang et al., 2022; Luo et al., 2019). The second line of works proposes dynamic embedding sizes for each feature (Liu et al., 2020, 2021; Zhao et al., 2021b; Qu et al., 2022). The third kind prevents recommender models from enumerating all high-order feature interactions when learning the interactions between features (Su et al., 2021, 2022). Model architecture search models explores different network architectures and determines the optimal architecture (Song et al., 2020; Cheng et al., 2022). Other components search focuses on optimize the loss function and feature interaction function (Zhao et al., 2021a; Meng et al., 2021).

### v5

@@ Abstract.
Latent factor models are the most popular backbones for today’s recommender systems owing to their prominent performance. Latent factor models represent users and items as real-valued embedding vectors for pairwise similarity computation, and all embeddings are traditionally restricted to a uniform size that is relatively large (e.g., 256-dimensional). With the exponentially expanding user base and item catalog in contemporary e commerce, this design is admittedly becoming memory-inefficient. To facilitate lightweight recommendation, reinforcement learning (RL) has recently opened up opportunities for identifying varying embedding sizes for different users/items. However, challenged by search efficiency and learning an optimal RL policy, existing RL-based methods are restricted to highly discrete, predefined embedding size choices. This leads to a largely overlooked potential of introducing finer granularity into embedding sizes to obtain better recommendation effectiveness under a given memory budget. In this paper, we propose continuous input embedding size search (CIESS), a novel RL-based method that operates on a continuous search space with arbitrary embedding sizes to choose from. In CIESS, we further present an innovative random walk-based exploration strategy to allow the RL policy to efficiently explore more candidate embedding sizes and converge to a better decision. CIESS is also model-agnostic and hence generalizable to a variety of latent factor recommender systems, whilst experiments on two real-world datasets have shown state-of-the-art performance of CIESS under different memory budgets when paired with three popular recommendation models. Code is available at https://github.com/qykcq/Continuous-Input-Embedding-Size-Search-For-Recommender-Systems.
@@ 1. Introduction
The root cause is that, each user/item embedding in the conventional embedding table shares the same fixed embedding size. An example from (Lian et al., 2020) shows that a recommender model embedding 10 million items into 256-dimensional vectors can exceed 9 GB memory consumption in a double-precision float system. As such, researchers have developed better solutions to compress the embedding table while maintaining its expressiveness. The most primitive recommender uses binary codes as an alternative to embeddings (Zhang et al., 2016), which was soon supplanted by another line of methods that compress fixed-sized embeddings into a lightly parameterized component. For example, codebooks have been proposed to store the embedding latent vectors efficiently in (Lian et al., 2020; Shi et al., 2020; Xia et al., 2023). (Xia et al., 2022) introduced the semi-tensor product operation to tensor-train decomposition to derive an ultra-compact embedding table, and (Kang et al., 2021) encoded users and items into hash codes and applied neural layers to learn their dense representations. However, these methods must be redesigned and retrained for different memory budgets to maintain optimal recommendation performance.
@@ 2. Methodology
CIESS has two main components that work alternately during training: (1) a recommendation model $F_{\Theta}(\cdot)$ parameterized by $\Theta$; and (2) the RL-based search function $G_{\Phi}(\cdot)$ parameterized by $\Phi$. A schematic view of CIESS’s workflow is shown in Figure 1. In each optimization iteration of CIESS, the recommender $F_{\Theta}(\cdot)$ adjusts its user/item embedding sizes to the ones provided by the policy $G_{\Phi}(\cdot)$, then updates its parameters $\Theta$ with training samples. Afterwards, $F_{\Theta}(\cdot)$ is evaluated on a hold-out dataset, where the top-$k$ recommendation quality can be measured by common metrics such as Recall at Rank $k$ (Recall@$k$) and Normalized Discounted Cumulative Gain at Rank $k$ (NDCG@$k$). Based on the recommendation quality, the search function $G_{\Phi}(\cdot)$ will be revised, and then updates its embedding size selection for each user/item for the next iteration. In what follows, we unfold the design of CIESS.
@@ 2.1. Base Recommender with Masked Embeddings
Let $\mathcal{U}$ and $\mathcal{V}$ be a set of users $u$ and items $v$, respectively. Their embedding vectors are stored in a real-valued embedding table $\mathbf{E}$ with the dimensionality of $(|\mathcal{U}|+|\mathcal{V}|)\times d_{max}$. It can be viewed as the vertical concatenation of all user and item embeddings $\mathbf{e}_{u_{1}};...;\mathbf{e}_{u_{|\mathcal{U}|}};\mathbf{e}_{v_{1}};...;\mathbf{e}_{v_{|\mathcal{V}|}}]$, where $d_{max}$ is the initial embedding dimension of all users/items in the full-size embedding table. In other words, $d_{max}$ is also the maximum embedding size in the search space.
@@
With the mask $\mathbf{M}$, for each user/item, we can retain the first $d_{n}$ elements of its full embedding while setting all succeeding dimensions to 0. It is worth noting that, performing embedding sparsification by masking unwanted dimensions with zeros is a commonly adopted approach in lightweight recommender systems (Liu et al., 2021; Lyu et al., 2022), as the resulting embedding table can take advantage of the latest sparse matrix storage techniques (Virtanen et al., 2020; Sedaghati et al., 2015) that bring negligible cost for storing zero-valued entries.
@@
where our target sparsity ratio $c$ ($0<c<1$) specifies the percentage of parameters to be pruned (i.e., zeroed out in our case) in the final sparse embedding matrix $\mathbf{E}\odot\mathbf{M}$.
@@ 2.2.2. State
where $f_{n}$ is the popularity of the user/item, normalized by the corresponding maximum/minimum frequency among all users $f_{max}=\max_{u\in\mathcal{U}}(f_{u})$$f_{min}=\min_{u\in\mathcal{U}}(f_{u})$ or items $f_{max}=\max_{v\in\mathcal{V}}(f_{v})$$f_{min}=\min_{v\in\mathcal{V}}(f_{v})$, which are observed from training data. $d_{n}$ is the current embedding size allocated to a user/item, and $q_{n}$ quantifies the changes in the recommendation quality when the embedding size decreases from $d_{max}$ to the current $d_{n}$. Compared with (Zhao et al., 2021b), incorporating this quality indicator into the state is able to help trace the impact from the most recent action (i.e., embedding size) to the recommendation effectiveness, which can encourage the policy network to better balance the memory cost and performance with the embedding sizes selected from a vast, continuous action space.
@@
[…]
@@ 2.2.3. Reward
The reward $r$ is the feedback to the current policy to guide subsequent embedding size adjustments. Given our goal stated in Eq.(5), the reward should reflect a blend of recommendation quality and space complexity. Therefore, we design the following pointwise reward for a user/item on their current embedding size $d_{u}$$d_{v}$ selected:
@@ 2.2.4. Action
At every iteration $i$, the policy network in the search function $G_{\Phi}$ predicts an action, i.e., dimension $d_{n}^{i}\in\mathbb{N}_{>0}$ from interval $1,d_{max}]$ given a user/item state $s_{n}^{i}$. The action is the embedding size for the corresponding user/item. The recommender $F_{\Theta}(\cdot)$ takes this action by altering the embedding sizes of users/items (i.e., updating $\mathbf{M}$ in practice), and yields a reward $r^{i}_{n}$ along with a subsequent state $s_{n}^{i+1}$. The tuple $(s^{i}_{n},d^{i}_{n},r^{i}_{n},s^{i+1}_{n})$ is defined as a transition, which are stored in a replay buffer $\mathcal{B}$ for subsequent training.
@@ 2.2.5. Actor and Critic
Then, starting from $\hat{d}^{i}_{n}$, we perform random walk (with replacement) to obtain a series of actions $\mathcal{Z}_{\hat{d}^{i}_{n}}$, where we specify $|\mathcal{Z}_{\hat{d}^{i}_{n}}|$ to be relatively small to balance efficiency and exploration thoroughness. After obtaining $\mathcal{Z}_{\hat{d}^{i}_{n}}$, we evaluate each action in $\mathcal{Z}_{\hat{d}^{i}_{n}}$ with the critic network $Q_{\mathcal{U}}(\cdot)$$Q_{\mathcal{V}}(\cdot)$ and greedily select the final action $d^{i}_{n}$ for iteration $i$:
@@ 2.3. Selective Retraining for Sparsified Embeddings
6:   /* Each iteration applies to all $n\in\mathcal{U}\cup\mathcal{V}$ */
@@
(15) $\forall\mathbf{M}\in\mathcal{M}_{c}:\;\;\;\;1-\frac{\|\mathbf{M}\|_{1,1}}{(|\mathcal{U}|+|\mathcal{V}|)\times d_{max}}\geq c$
@@ 4. Related Work
4.1. Latent Factor Recommenders
Neural networks have exhibited superior ability in solving recommendation tasks. Since MLP can learn the non-linear interactions between users and items, feature representation learning with MLP has been commonplace. He et al. (He et al., 2017) proposed Neural Collaborative Filtering consisting of an MLP component that learns the non-linear user-item interactions and a generalized matrix factorization component that generalizes the traditional matrix factorization. Cheng et al. (Cheng et al., 2016) proposed Wide&Deep connecting a wide linear model with a deep neural network for the benefits of memorization and generalization. Apart from MLP-based models, graph-based methods (Wang et al., 2019; He et al., 2020; Yu et al., 2022) have also demonstrated promising capability. Wang et al. (Wang et al., 2019) proposed the graph neural network-based model NGCF to model the user-item interactions with a bipartite graph and propagate user-item embeddings on the embedding propagation layers. Similarly, He et al. (He et al., 2020) resorted to Graph Convolution Network (GCN) and proposed LightGCN that simplified NGCF by including only its neighborhood aggregation component for collaborative filtering. LightGCN performs neighborhood filtering by computing the weighted sum of the user and item embeddings of each layer. Another promising direction, Factorization Machine (Rendle, 2010), along with its deep variants QNFM (Chen et al., 2021a), DeepFM (Guo et al., 2017) and xLightFM (Jiang et al., 2021) have been studied. (Blondel et al., 2016) was also created to mine higher-order interactions. In addition, heterogeneous data such as textual (Gong and Zhang, 2016; Li et al., 2017; Chen et al., 2020) and visual (Wang et al., 2017; Lee et al., 2018) data has also been exploited by several CNN or RNN-based models. Most of these recommenders utilize vectorized embeddings, which can be optimized by CIESS for memory-efficient embeddings.
@@ 4.1. Latent Factor Recommenders
4.2. AutoML for Recommendation.
Designing deep recommender models heavily relies on the expertise of professionals. To alleviate the need for human engagement, Automated Machine Learning for Recommender Systems has been studied to automate the design process for recommender models in a task-specific and data-driven manner. So far several research directions including feature selection search (Wang et al., 2022; Luo et al., 2019), embedding dimension search (Liu et al., 2021, 2020; Zhao et al., 2021b; Qu et al., 2022), feature interaction search (Su et al., 2021, 2022), model architecture search (Song et al., 2020; Cheng et al., 2022), and other components search (Zhao et al., 2021a; Meng et al., 2021) have been proposed (Zheng et al., 2022). The first kind reduces computation cost by filtering feature fields based on learnable importance scores (Wang et al., 2022; Luo et al., 2019). The second line of works proposes dynamic embedding sizes for each feature (Liu et al., 2020, 2021; Zhao et al., 2021b; Qu et al., 2022). The third kind prevents recommender models from enumerating all high-order feature interactions when learning the interactions between features (Su et al., 2021, 2022). Model architecture search models explores different network architectures and determines the optimal architecture (Song et al., 2020; Cheng et al., 2022). Other components search focuses on optimize the loss function and feature interaction function (Zhao et al., 2021a; Meng et al., 2021).